How do I set up the equation for "A bank teller cashed a check for $1010 using $20 bills and $50 bills. The teller handed 26 more $20 bills than $50 bills to the customer. Determine how many of each denomination were handed to the customer."?
Answer by nerdybill(7384)

How do I set up the equation for "A bank teller cashed a check for $1010 using $20 bills and $50 bills. The teller handed 26 more $20 bills than $50 bills to the customer. Determine how many of each denomination were handed to the customer."?
Let x = number of $50 bills
then
x+26 = number of $20 bills
.
Equation:
50x + 20(x+26) = 1010
50x + 20x+520 = 1010
70x+520 = 1010
70x = 490
x = 70 (number of $50 bills)
.
$20 bills:
x+26 = 70+26 = 96